Find the discriminant of the equation and give the number and type of solution of the equation. x^2+6x-16=0
Answer by venugopalramana(3286)

Find the discriminant of the equation and give the number and type of solution of the equation. x^2+6x-16=0
FOR
AX^2+BX+C=0.....
D=B^2-4AC
D = 6^2-4*1*(-16)=36+64=100....
IT IS +VE........HENCE ROOTS ARE REAL AND DISTINCT
IT IS A PERFECT SQUARE.HENCE ROOTS ARE RATIONAL